OpenStudy (anonymous):

Solve. x2 + 8x = –13

OpenStudy (anonymous):

(x+4)^2-3=0 x^2+8x+13=0

OpenStudy (anonymous):

I'm still confused

OpenStudy (anonymous):

What you need help in?

OpenStudy (anonymous):

Write out x(x+8)=-13

OpenStudy (anonymous):

The final answer?

OpenStudy (anonymous):

Final answer is x^2+8x+13=0

OpenStudy (anonymous):

Add '-13' to each side of the equation. 13 + 8x + -13 + x2 = 0 + -13 Reorder the terms: 13 + -13 + 8x + x2 = 0 + -13 Combine like terms: 13 + -13 = 0 0 + 8x + x2 = 0 + -13 8x + x2 = 0 + -13 Combine like terms: 0 + -13 = -13 8x + x2 = -13 The x term is 8x. Take half its coefficient (4). Square it (16) and add it to both sides. Add '16' to each side of the equation. 8x + 16 + x2 = -13 + 16 Reorder the terms: 16 + 8x + x2 = -13 + 16 Combine like terms: -13 + 16 = 3 16 + 8x + x2 = 3 Factor a perfect square on the left side: (x + 4)(x + 4) = 3 (x + 4) equal to 1.732050808 and -1.732050808

OpenStudy (anonymous):

here are the choices –4 ± √3 –4 ± √17 –4 ± √21 –4 ± √29

OpenStudy (anonymous):

Its the first one

OpenStudy (anonymous):

–4 ± √29

OpenStudy (anonymous):

Wait what? which one is the real correct answer?

OpenStudy (anonymous):

Omg everyone is confusing her

OpenStudy (anonymous):

–4 ± √29 enter it in your calculator as -4+√29

OpenStudy (anonymous):

Yeah, but you both told me different answers I'm super confused

OpenStudy (anonymous):

The other person does not know what they are talking about

OpenStudy (anonymous):

How did you get the answer for it then?

OpenStudy (anonymous):

SImpleton must be smart, Well sometimes guesses work =)

OpenStudy (anonymous):

i did x + 4 = 1.732050808 Simplifying x + 4 = 1.732050808 Reorder the terms: 4 + x = 1.732050808 Solving 4 + x = 1.732050808 Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Add '-4' to each side of the equation. 4 + -4 + x = 1.732050808 + -4 Combine like terms: 4 + -4 = 0 0 + x = 1.732050808 + -4 x = 1.732050808 + -4 Combine like terms: 1.732050808 + -4 = -2.267949192 x = -2.267949192 Simplifying x = -2.267949192 that gives you -4-√29